OpenStudy (anonymous):

64^(3 – x) = 4^2x solve and explain how

OpenStudy (lgbasallote):

hint: \(\large 64 = 4^3\)

OpenStudy (anonymous):

idk

OpenStudy (lgbasallote):

replace 64 with 4^3 in your equation what will you have?

OpenStudy (anonymous):

4^3-x=4^2x???

OpenStudy (lgbasallote):

\[\LARGE (4^3)^{(3-x)} = 4^{2x}\] got it?

OpenStudy (anonymous):

ok yeah

OpenStudy (lgbasallote):

okay...then remember that \[\LARGE (a^b)^c = a^{bc}?\]

Parth (parthkohli):

Now, make up an equation of the exponents as the bases are same.

OpenStudy (anonymous):

4^(9-3x)=4^2x?

Parth (parthkohli):

Yes. \(\Large \color{MidnightBlue}{\Rightarrow 9 - 3x = 2x }\)

OpenStudy (anonymous):

yea

OpenStudy (lgbasallote):

right! another property.. \[\LARGE a^x = a^y \rightarrow x = y\] that meant that if two similar bases with different exponents are equal then their exponents are equal...capsiche?

OpenStudy (anonymous):

yea

Parth (parthkohli):

^ I like this lgba :P

OpenStudy (anonymous):

9=5x

OpenStudy (anonymous):

x=9/5?

Parth (parthkohli):

Divide both sides by 5 now.

Parth (parthkohli):

Yes.

OpenStudy (lgbasallote):

\[\Huge \color{red}{\checkmark}\]

OpenStudy (lgbasallote):

lol im a fail

OpenStudy (anonymous):

y?

OpenStudy (lgbasallote):

i wrote the latex for that check wrong the first time haha